# Artificial Intelligence in Creative Industries: Advances Prior to 2025

### v2

@@ 2.2 Large language models
LLMs are based on transformer models with self-attention mechanisms as core modules. Training LLMs comprises two steps: i) pre-training in unsupervised learning manner, and ii) fine-tuning to a specific task or prompt-tuning (or prompt engineering) for better user inputs. The models are first ‘pre-trained’ with a large amount of unlabelled text data to learn the meaning of words, and the relationships between words before using it to adapt to a downstream task. This is why OpenAI calls their models generative pre-trained transformers (GPT).
Fine-tuning involves training the model on new datasets. The drawback is that these data need to be large enough to generalize a new task. Prompt-tuning and prompt engineering are relatively new disciplines for developing and optimizing prompts to efficiently use language models. Prompts is a component that guides the way AI models interpret and respond to user queries. Prompt engineering is the process of structuring text or phrasing that guide the model towards generating the desired output. This relies heavily on trial and error, and an understanding of how the model responds. Prompt-tuning, on the other hand, involves training a small set of parameters before utilizing the LLM, thus requiring a relatively small amount of new data. This approach essentially converts text inputs into task-specific virtual inputs, referred to as tokens, while the pre-trained LLM remains unchanged [27]. As it is AI-based design, the main drawback of prompt-tuning is lack of interpretability. This paradigm has however extended to other domains, such as visual prompt tuning [28]. For a comprehensive survey of LLMs, please refer to [29].
To date, there are many LLM platforms as Fig. 2 (a) shows their timeline. Many publications on surveys and evaluations of LLMs are available [29, 30, 31]. This includes FLASK (Fine-grained Language Model Evaluation based on Alignment SKill Sets) [32] evaluating LLMs based on 12 fine-grained skills for comprehensive language model evaluation: logical correctness, logical robustness, logical efficiency, factuality, commonsense understanding, comprehension, insightfulness, completeness, metacognition, conciseness, readability, and harmlessness. Their evaluation results are shown in Fig. 2 (b).
@@ 2.3 Diffusion Models
A generative model, in the context of generative AI, exploits machine learning to learn a probability distribution of the training data to generate new data samples. The very first models are based on Autoencoders (AEs) that learn to encode input data into a lower-dimensional representation (latent space) and then decode it back to its original form. The specific type of AEs, Variational Autoencoders (VAEs) [33], learns the latent space as statistic parameters of probabilistic distributions, leading to significant improvement of the generated results. Concurrently, in the same year, Goodfellow et al. [34] introduced an alternative architecture known as a Generative Adversarial Network (GAN). GANs comprise two competing AI modules: the generator, which creates a sample, and the discriminator, which determines whether the received sample is real or generated. When comparing VAEs to GANs, VAEs exhibit greater stability during training, whereas GANs excel at producing realistic images. More details about AEs and GANs for creative technologies can be found in our previous review [8].
@@
The important factor driving the rapid growth of generative AI is the development of diffusion probabilistic models, or in short, diffusion models (DMs). The first DM was introduced in 2015 by Sohl-Dickstein et al. [35], using Nonequilibrium Thermodynamics. However, it took 5 years for DMs to generate desirable results. The era of DMs began with Denoising Diffusion Probabilistic Models (DDPMs) proposed by Ho et al. [36] in 2020 and Score-based diffusion models proposed by Song et al. [37] in 2021. These are a simplified process using a denoising autoencoder to approximate Bayesian inference. In brief, the models leverage a diffusion process to learn a probability distribution of the input data. As the name suggests, the data is diffused by gradually adding noise at each iteration step as shown in Fig. 1 (b). A deep neural network (DNN) is then trained to remove this noise, called the denoising process or reverse process. Consequently, the trained model uses random noise to generate data with characteristics similar to those of the training samples. Comparing to GANs, the DMs provide higher diversity samples [38] and the training process is much more stable and do not suffer from mode collapse; however, the DMs are computationally intensive and require longer training times compared to GANs. The complexity is significantly reduced by training the DMs in latent space. Latent Diffusion Models (LDM) [26] use pretrained networks to convert images to feature maps, and perform training on a low-dimensional space. The diagram of LDM is shown in Fig. 1 (c).
@@
Generating a synthesized sample at random might not be particularly useful, especially for creative industries. Therefore, conditional diffusion models have been proposed, with a wide range of applications such as text-to-sound, text-to-images, and image-to-videos. For DMs, the conditional distributions are modelled using a conditional denoising autoencoder. Classifier guidance was introduced in [38] to improve a diffusion generator in generating images of a desired class. For example, when we provide the model with more information, such as ‘a flower’, the DM will synthesize a variety of flower images, as the word ‘flower’ guides the model toward the latent distribution that is formed by various images of flowers. The work in [39] simply refines the latent space of well-trained unconditional DDPM so that the higher-level semantics of the synthetic samples are similar to the reference (conditioning). The LDM [26] offers more flexible conditional image generators by adding cross-attention layer (referred to Transformers in Section 2.1) to the denoising autoencoder. A survey on the methods and applications of DMs prior to 2024 can be found in [40].
@@ 2.4 Implicit Neural Representations
Implicit Neural Representations (INR), also called neural fields, neural implicits, and coordinate-based neural networks, represent input content implicitly through learned functions $F$, as shown in Eq. 3. They can be seen as fields $x$ (represented by a scalar, vector, or tensor with a value, such as magnetic field in physics) that are fully or partially parameterized by a neural network $\Phi$, typically an MLP [41].
@@
The concept might sound complex, but the process is actually very straightforward. For example, in the case of an image, the coordinate of each pixel $(x,y)$ contains colors $(r,g,b)$. The INR inputs $(x,y)$ to the MLP and learns to give the output $(r,g,b)$. The weights and biases of the MLP now represent such an image. Usually, the number of parameters of the MLP is smaller than the total number of pixels multiplied by 3, accounting for the 3 color channels. Hence, one of its applications is data compression [42]. Moreover, the INR can handle complex and high-dimensional data efficiently, bringing attention in visual computing, such as 3D scene reconstruction.
Traditional MLPs employ ReLU (rectified linear unit) for non-linear activation due to its simplicity. However, Sitzmann et al. [43] demonstrate that using periodic functions, such as sinusoids, is more suitable for representing complex natural signals, offering a better fit to the first- and second-order derivatives of the signals. However, this activation could cause ringing artifacts. Saragadam et al. instead proposed using complex Gabor wavelets [44], which learn to represent high frequencies better and simultaneously are robust to noise.
One of the fastest-growing areas that exploits INRs is Neural Radiance Fields (NeRF), evidenced by 57 papers presented at CVPR, the largest annual conference in computer vision, in 2022 and which grew to 175 papers in 2023777https://markboss.me/post/nerf_at_cvpr23/. First introduced in 2020 by Mildenhall et al. [45], NeRF is a form of neural rendering, a subset of generative AI, that generates novel views of a scene based on a partial set of 2D images. It achieves this by learning a mapping from 3D spatial coordinates and view directions $(x,y,z,\theta,\phi)$ to colors and density $(r,g,b,\sigma)$. This implicit representation allows NeRF to handle complex scenes with varying geometry and appearance, resulting in highly realistic renderings that include accurate lighting, shadows, and reflections. More detail can be found in Section 3.5.2.
@@
audio/music [48, 49] [50, 51]
Image [48, 52] [26, 53, 54, 55, 56, 52, 57]
Animation/video [58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68] [69, 70, 71, 72, 62, 66, 68]
@@
Information Text categorization [80, 81, 82]
@@
and Post Super-resolution [111, 112, 113, 114, 115, 116, 99, 117] [118, 119, 120] [121, 118, 122, 120, 123]
Production Restoration [124, 125, 115, 126, 99, 127, 111, 18, 128, 5, 129, 130, 131, 132, 133] [102, 122, 134, 135, 136] [137]
@@
Information Segmentation [149, 150, 151, 5, 152, 153, 154, 155] [156, 157, 158] [159, 160]
Extraction Recognition [161, 11, 162, 16, 19, 17, 163, 154, 164] [165, 166]
@@
1 Trans./Attn. include both transformers and CNN-based architectures that use attention module.
@@ 3 Advanced AI for the creative industries
Similar to our previous paper reviewing AI for the creative industries, we categorize the applications and the corresponding AI-based solutions as shown in Table 1. However, the subsections may differ as generative AI affects each application differently.
@@ 3.1 Content creation
Content creation stands as a fundamental activity of artists and designers and the term ‘AI art’ refers to either created with the assistance of AI algorithms or entirely by AI systems. It is in various digital forms, including images, texts, audio, and videos. The roots of AI art can be traced back to before the 2000s, exemplified by AARON, a computer program initiated in 1972 to autonomously produce paintings and drawings [231]. The practicality of AI art has actually begun with advancements in deep learning, particularly GAN from 2014 and, more recently, transformers, DMs and INRs. Note that when referring to real camera content acquisition using AI, commonly known as intelligent cinematography, we direct the reader to the review in [232].
@@ 3.1.1 Text generation, script and journalism
In the era of LLMs, AI writing tools have been widely used to help with various writing tasks, including generate written content of articles, blog posts, essays, and reports. These tools go beyond mere grammar and spelling checks; they boast advancements enabling them to analyze the style and tone of written material, adding images, videos and tables, offering suggestions to enhance clarity, coherence, and overall readability [233]. Moreover, AI tools extend their utility beyond content generation by automating tasks like keywords, meta tags, and descriptions, thereby increasing search rankings using search engine optimization (SEO). Additionally, they support the process of publishing across multiple online platforms. Transformers are used to generate image captions by combining information of the images with a word prefix or questions [46].
@@
As discussed earlier in this paper, chatbots are now powered by LLMs, effectively simulating human conversation. These fundamental LLMs are specialized for specific tasks, for instance, journalist AI and blog AI writers [^8] For example, see https://tryjournalist.com/, which generate content with layouts suitable for print or online publication. Additionally, there are AI tools designed to detect AI-generated content (e.g., for checking for copyrights), AI-writing styles and content originality to ensure the natural sound of articles. Undoubtedly, generative AI is reshaping the way artists and journalists operate. For an in-depth exploration of the impact and implications of these technological advancements on news organizations, refer to the survey conducted by Beckett et al. [236].
Generating text and scripts automatically can also be done through image and video inputs without text prompts (e.g., image captioning [237]) and with text prompts. The technical term of these approaches is called Vision Language Models (VLMs), which are multimodal models that learn from images and text. The most common and prominent models often consist of an image encoder, an embedding projector to align image and text representation, often via a dense neural network, and a text decoder stacked in this order. The most well-known technique is Contrastive Language-Image Pre-training (CLIP) [238]. More recent work in [47] scales up the vision vocabulary by incorporating new image features into the existing CLIP model, resulting in improved content understanding. The comprehensive survey of VLMs for vision tasks can be found in [239].
@@ 3.1.2 Audio and Music generation
Similar to language models, AI-based music generation has rapidly advanced due to unsupervised learning on large datasets and the use of transformers (see Section 2.2). Examples of such software include MuseNet [^9] https://openai.com/research/musenet, Magenta Studio [^10] https://magenta.tensorflow.org/studio, and Musicfy [^11] https://musicfy.lol/. These tools can assist in composing music by learning complex musical patterns, predicting the next word or music note in a sequence, and mixing specified instruments. Moreover, the AI tools can convert one type of sound into another, such as from whistling to violin or from flute to saxophone [^12] See an example by Ummet Ozcan at https://www.youtube.com/watch?v=lI1LCfTx2lI. This capability is invaluable for artists who may not be proficient in playing all the instruments they wish to incorporate, saving both time and costs. March 2024, Suno has released their first model capable of producing radio-quality music that can be created in 2 minutes [^13] https://www.suno.ai/blog/v3. Later, Udio  [^14] https://www.udio.com/, was launched. It offers a prompt to create lyrics and music with a maximum duration of 90 seconds, and also appears to have at least some awareness of copyright.
@@
AI voice software changes voice from one person to another and enable users to train the model to convert other people’s voices into their own, e.g. lalals [^15] https://lalals.com/, Kits [^16] https://www.kits.ai/, Media.io [^17] https://www.media.io/online-voice-changer.html, etc. Certain software, such as Voice.ai [^18] https://voice.ai/, even offer real-time voice changing capabilities. The technologies behind this use the transformer to learn voice features and patterns in mel-spectrogram. For example, the framework proposed in [51] uses a method based on DM with transformer backbone to turn text input into a mel-spectrogram using the vector quantized variational autoencoder (VQ-VAE) [240]. Next, this mel-spectrogram is transformed into a sound wave. Unlike a regular spectrogram, the mel-spectrogram is based on the mel-frequency scale, which offers higher resolution for lower frequencies. Voice style transfer often use zero-shot learning (a model is trained to recognize classes or categories it has never seen during training) [49] or few-shot learning (a model is trained with only one or a few examples per class) [241].
@@
Another emerging AI technology is in the field of spatial audio. In 2022, Apple Music revealed that, in just over a year, more than 80% of its worldwide subscribers were enjoying the spatial audio experience, with monthly plays in spatial audio increasing by over 1,000% [^19] https://www.apple.com/uk/newsroom/2023/01/apple-celebrates-a-groundbreaking-year-in-entertainment/. With head tracking, this technology significantly enhances the immersive experience. Masterchannel has launched SpatialAI [^20] https://platform.masterchannel.ai/spatial, claiming it to be the world’s first spatial mastering AI. It processes audio files and returns an optimized track for streaming platforms, along with an individually optimized stereo version for traditional distribution. All these advancements leverage transformer-based technologies.

### v3

@@ 2.2 Large language models
LLMs are based on transformer models using self-attention mechanisms as their core modules. Training LLMs comprises two steps: i) pre-training in an unsupervised learning manner, and ii) fine-tuning to a specific task or prompt-tuning for better user inputs. The models are first ‘pre-trained’ with a large amount of unlabelled text data to learn the meaning of words, and the relationships between those words, before using it to adapt to a downstream task. This is why OpenAI refers to their model as a Generative Pre-trained Transformer (GPT).
Fine-tuning involves training the model on new datasets. The drawback is however that these data need to be large enough to ensure generalization to new tasks. Prompt-tuning and prompt engineering are relatively new disciplines for developing and optimizing prompts to efficiently use language models. Prompts guide the way AI models interpret and respond to user queries. Prompt engineering is the process of structuring text or phrasing that guide the model towards generating the desired output. This relies heavily on trial and error, and an understanding of how the model responds. Prompt-tuning, on the other hand, involves training a small set of parameters before utilizing the LLM, thus requiring a relatively small amount of new data. This approach essentially converts text inputs into task-specific virtual inputs, referred to as tokens, while the pre-trained LLM remains unchanged [25]. The main drawback of prompt-tuning is lack of interpretability. This paradigm has however extended to other domains, such as visual prompt tuning [26]. For a comprehensive survey of LLMs, please refer to [27].
To date, there are many LLM platforms as shown in Fig. 2. Fig. 2(a) shows their timeline. Many surveys and evaluations of LLMs are also available [27, 28, 29]. These include FLASK (Fine-grained Language Model Evaluation based on Alignment SKill Sets) [30] which evaluates LLMs based on 12 fine-grained skills for comprehensive language model evaluation: logical correctness, logical robustness, logical efficiency, factuality, commonsense understanding, comprehension, insightfulness, completeness, metacognition, conciseness, readability, and harmlessness. Evaluation results from FLASK are shown in Fig. 2 (b).
@@ 2.3 Diffusion Models
A generative model, in the context of AI, exploits machine learning to learn a probability distribution of the training data to generate new data samples. The very first models were based on Autoencoders (AEs) that learn to encode input data into a lower-dimensional representation (latent space) and then decode it back to its original form. A specific type of AE, a variational autoencoders (VAE) [31], learns the latent space as statistical parameters of probabilistic distributions, leading to significant improvement of the generated results. Concurrently, Goodfellow et al. [32] introduced an alternative architecture known as a Generative Adversarial Network (GAN). GANs comprise two competing AI modules: a generator, which creates a sample, and a discriminator, which determines whether the received sample is real or generated. When comparing VAEs to GANs, VAEs exhibit greater stability during training, whereas GANs excel at producing realistic images. More details about AEs and GANs for creative technologies can be found in our previous review [8].
@@
An important factor in driving the rapid growth of generative AI has been the development of diffusion probabilistic models (referred to as diffusion models (DMs)). The first DM was introduced in 2015 by Sohl-Dickstein et al. [33], using Nonequilibrium Thermodynamics. However, it took a further 5 years for DMs to generate desirable results: the era of DMs began with Denoising Diffusion Probabilistic Models (DDPMs) proposed by Ho et al. [34] in 2020 and Score-based diffusion models proposed by Song et al. [35] in 2021. These involve a simplified process using a denoising autoencoder to approximate Bayesian inference. In brief, the models leverage a diffusion process to learn a probability distribution of the input data. As the name suggests, the data is diffused by gradually adding noise at each iteration step as shown in Fig. 1 (b). A deep neural network (DNN) is then trained to remove this noise, called the denoising process or reverse process. Consequently, the trained model uses random noise to generate data with characteristics similar to those of the training samples. Comparing to GANs, DMs provide higher diversity samples [36] and a training process that is much more stable and does not suffer from mode collapse. DMs are however computationally intensive and require longer training times compared to GANs. The complexity can significantly reduced by training the DMs in latent space. Latent Diffusion Models (LDM) [24] use pretrained networks to convert images to feature maps, and perform training on a low-dimensional space. The diagram of LDM is shown in Fig. 1 (c).
@@
Generating a synthesized sample at random might not be particularly useful, especially for creative industry applications. Therefore, conditional diffusion models have been proposed, supporting a wide range of applications such as text-to-sound, text-to-images, and image-to-videos. For DMs, the conditional distributions are modelled using a conditional denoising autoencoder. Classifier guidance was introduced in [36] to improve the generation of images of a desired class. For example, when we provide the model with information, such as ‘a flower’, the DM will synthesize a variety of flower images, as the word ‘flower’ guides the model toward the latent distribution that is formed by various images of flowers. The work in [37] simply refines the latent space of well-trained unconditional DDPM so that the higher-level semantics of the synthetic samples are similar to the reference (conditioning). The LDM [24] offers more flexible conditional image generators by adding cross-attention layer (referred to Transformers in Section 2.1) to the denoising autoencoder. A survey on the methods and applications of DMs prior to 2024 can be found in [38].
@@ 2.4 Implicit Neural Representations
Implicit Neural Representations (INR), also called neural fields, neural implicits or coordinate-based neural networks, represent input content implicitly through learned functions $F$, as shown in Eq. 3. They can be considered as fields $x$ (represented by a scalar, vector, or a tensor with a value, such as magnetic field in physics) that are fully or partially parameterized by a neural network $\Phi$, typically an MLP [39].
@@
Although this concept appears complex, the process is actually very straightforward. For example, in the case of an image, the coordinates of each pixel $(x,y)$ contain color information $(r,g,b)$. The INR inputs $(x,y)$ to the MLP and learns to provide the output $(r,g,b)$. The weights and biases of the MLP now represent such an image. Usually, the number of parameters of the MLP is smaller than the total number of pixels multiplied by 3, accounting for the 3 color channels. Hence, one of its emerging applications is in data compression [40]. Moreover, the INR can handle complex and high-dimensional data efficiently, attracting attention for visual computing applications such as 3D scene reconstruction.
Traditional MLPs employ ReLU (rectified linear unit) for non-linear activation due to its simplicity. However, Sitzmann et al. [41] demonstrated that using periodic functions, such as sinusoids, are more suitable for representing complex natural signals, offering a better fit to the first- and second-order derivatives of the signals. However, this activation can cause ringing artifacts. Saragadam et al. instead proposed using complex Gabor wavelets [42], which learn to represent high frequencies better and simultaneously are robust to noise.
One of the fastest-growing areas that exploits INRs is Neural Radiance Fields (NeRF), evidenced by 57 papers presented at CVPR, the largest annual conference in computer vision, in 2022 growing to 175 papers in 2023101010https://markboss.me/post/nerf_at_cvpr23/, before dropping to 71 in 2024, largely due to competition from 3D Gaussian Splatting [^11] https://github.com/Yubel426/NeRF-3DGS-at-CVPR-2024. First introduced in 2020 by Mildenhall et al. [43], NeRF is a form of neural rendering, a subset of generative AI, that generates novel views of a scene based on a partial set of 2D images. It achieves this by learning a mapping from 3D spatial coordinates and view directions $(x,y,z,\theta,\phi)$ to colors and density $(r,g,b,\sigma)$. This implicit representation allows NeRF to handle complex scenes with varying geometry and appearance, resulting in highly realistic renderings that include accurate lighting, shadows, and reflections. More detail can be found in Section 3.5.2.
@@
audio/music [46, 47] [48, 49, 50]
Image [46, 51] [24, 52, 53, 54, 55, 51, 56, 57, 58]
Animation/video [59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69] [70, 71, 72, 73, 63, 67, 69, 74, 75]
@@
Information Text categorization [83, 84, 85, 86]
@@
and Post Super-resolution [115, 116, 117, 118, 119, 120, 103, 121, 122] [123, 124, 125, 126, 122] [127, 123, 128, 125, 129]
Production Restoration [130, 131, 119, 132, 103, 133, 115, 18, 134, 5, 135, 136, 137, 138, 139, 140, 141, 142, 141, 143] [106, 128, 144, 145, 146, 126, 147] [148]
@@
Information Segmentation [161, 162, 163, 5, 164, 165, 166, 167, 168] [169, 170, 171] [172, 173]
Extraction Recognition [174, 11, 175, 16, 176, 17, 177, 166, 178, 179, 180] [181, 182, 183, 184]
@@
1 Trans./Attn. include transformers, mamba and CNN-based architectures that use attention module.
@@ 3 Advanced AI for the creative industries
Similarly to our previous (2021) review of AI for the creative industries [8], Table 1 categorizes applications and corresponding AI-based solutions. These areas are explored in more detail below.
@@ 3.1 Content creation
Content creation is a fundamental activity of artists and designers and the term ‘AI art’ refers to artforms created with the assistance of an AI algorithms or entirely by an AI system. This can refer to various digital forms including images, texts, audio, and videos. The roots of AI art can be traced back to the 20th century, exemplified by AARON, a computer program initiated in 1972 to autonomously produce paintings and drawings [256]. The practicality of AI art has been enhanced with advancements in deep learning, particularly GANs from 2014 and, more recently, transformers, DMs and INRs.
@@ 3.1.1 Text generation, script and journalism
In the era of LLMs, AI writing tools have been widely used to assist various writing tasks, including generation written articles, blog posts, essays, and reports. These tools go beyond mere grammar and spelling checks; they boast advancements enabling them to analyze the style and tone of written material, adding images, videos and tables, offering suggestions to enhance clarity, coherence, and overall readability [257]. Moreover, AI tools extend their utility beyond content generation by automating tasks like keyword generation, meta tags, and descriptions, thereby increasing search rankings using search engine optimization (SEO). Additionally, they support the process of publishing across multiple online platforms. Transformers have been used to generate image captions by combining information from the images with a word prefix or questions [44].
@@
As discussed earlier, chatbots are now powered by LLMs, effectively simulating human conversation. These fundamental LLMs are specialized for specific tasks. For instance journalist AI and blog AI writers [^12] For example, see https://tryjournalist.com/ generate content with layouts suitable for print or online publication. Additionally, AI tools exist that are designed to detect AI-generated content (e.g., for checking for copyright), AI-writing styles, content originality and to ensure the naturalness and flow of articles. Undoubtedly, generative AI is reshaping the way artists and journalists operate. For an in-depth exploration of the impact and implications of these technological advancements on news organizations, refer to the survey conducted by Beckett et al. [261].
Generating text and scripts automatically can also be done through image and video inputs without text prompts (e.g., image captioning [262]) and with text prompts. These approaches are referred to as Vision Language Models (VLMs): multimodal models that learn from images and text. The most common and prominent models often consist of an image encoder, an embedding projector to align image and text representation, often via a dense neural network, and a text decoder stacked in this order. The most well-known technique is Contrastive Language-Image Pre-training (CLIP) [263]. More recent work in [45] scales up the vision vocabulary by incorporating new image features into the existing CLIP model, resulting in improved content understanding. A comprehensive survey of VLMs for vision tasks can be found in [264].
@@ 3.1.2 Audio and music generation
Similar to language models, AI-based music generation has rapidly advanced due to unsupervised learning on large datasets and the use of transformers (see Section 2.2). Examples of such systems include MuseNet [^13] https://openai.com/research/musenet, Magenta Studio [^14] https://magenta.tensorflow.org/studio, and Musicfy [^15] https://musicfy.lol/. These tools assist in music composition by learning complex musical patterns, predicting the next word or music note in a sequence, and mixing specified instruments. Moreover, AI tools can convert one type of sound into another, such as from whistling to violin or from flute to saxophone [^16] See an example by Ummet Ozcan at https://www.youtube.com/watch?v=lI1LCfTx2lI. This capability is invaluable for artists who may not be proficient in playing all the instruments they wish to incorporate, saving both time and costs. In 2024, Suno has released a model capable of producing radio-quality music that can be created in 2 minutes [^17] https://www.suno.ai/blog/v3. Later, Udio  [^18] https://www.udio.com/, was launched. This offers a prompt to create lyrics and music with a maximum duration of 90 seconds, and also appears to have, at least some, awareness of copyright.
@@
AI voice software changes vocalizations from one person to another, for example enabling users to train the model to convert other people’s voices into their own, e.g. lalals [^19] https://lalals.com/, Kits [^20] https://www.kits.ai/, Media.io [^21] https://www.media.io/online-voice-changer.html, etc. Certain software, such as Voice.ai [^22] https://voice.ai/, even offers real-time voice changing capabilities. The technologies behind this uses a transformer to learn voice features and patterns in mel-spectrogram form. For example, the framework proposed in [49] uses a DM-based method with a transformer backbone to turn text input into a mel-spectrogram using the vector quantized variational autoencoder (VQ-VAE) [265]. Next, this mel-spectrogram is transformed into a sound wave. Unlike a regular spectrogram, the mel-spectrogram is based on the mel-frequency scale, which offers higher resolution for lower frequencies. Voice style transfer often uses zero-shot learning (a model is trained to recognize classes or categories that it has never encountered during training) [47] or few-shot learning (a model trained with only one or a few examples per class) [266]. Stable Audio Open [50] introduces a text-conditioned generative model for non-speech audio, trained on Creative Commons licensed data, capable of producing state-of-the-art 44.1kHz stereo audio.
@@
Another emerging AI technology application is in the field of spatial audio. In 2022, Apple Music revealed that, in just over a year, more than 80% of its worldwide subscribers were enjoying the spatial audio experience, with monthly plays in spatial audio increasing by over 1,000% [^23] https://www.apple.com/uk/newsroom/2023/01/apple-celebrates-a-groundbreaking-year-in-entertainment/. With head tracking, this technology significantly enhances the immersive experience. Masterchannel has launched SpatialAI [^24] https://platform.masterchannel.ai/spatial, claiming it to be the world’s first spatial mastering AI. This processes audio files and returns an optimized track for streaming platforms, along with an individually optimized stereo version for traditional distribution. All these advancements leverage transformer-based technologies.